2. Two cyclist start biking from a trails start 3 hours apart. The second cyclist travels at 10 miles per hour and starts 3 hours after the first cyclist who is traveling at 6 miles per hour. How much time will pass before the second cyclist catches up with the first from the time the second cyclist started biking?
A. 2 hours
B. 4 1/2 hours
C. 5 3/4 hours
D. 6 hours
E. 7 1/2 hours
Please can you show me the solution.

Answer by nyc_function(2741)

The first cyclist has traveled 18 miles in three hours.
The second cyclist is gaining on him at 10 - 6 = 4mph and has to make up 18 miles at that 4mph.
18 / 4 = 9/2 = 4 1/2 (B) <------------------ Answer